A lagged copper rod of uniform cross sectional area has lenght of 60cm the free ends of the rod are maintained at 20°C and 0°C respectively at a steady state.calculate the temperature at a point 20cm from the high temperature end


1
Expert's answer
2021-02-11T08:54:05-0500

Let's consider the heat equation in a form:


"u_{xx}=u_{t}"

where "u(x,t)" is the temperature that depends on position "x" and time "t". Subscripts denote partial derivatives. At the steady state the temperature "u = T(x)" should not depend on the time, thus: "T_t = 0". Solving the remaining equation, find:


"T_{xx} = 0\\
T_x = C_1\\
T(x) = C_1x+C_2"

where "C_1,C_2" are some constants. Substituting the boundary value on the one end, obtain:


"T(0) = C_1\cdot 0+ C_2 =0\\
C_2 =0"


On another end:


"T(60) =C_1\cdot 60+0 =20\\
C_1 = 1/3"

Thus, the equation is:


"T(x) = \dfrac{x}{3}"

And the temperature at "x = 20cm" is:


"T(20) = \dfrac{20}{3} \approx 6.67\degree C"

Answer. "6.67\degree C".